# A Hitchhiker’s Guide to Privacy-Preserving Digital Payment Systems: A Survey on Anonymity, Confidentiality, and Auditability

### v2

@@ 4.3.5. Linkability and Unlinkability
  • Address reuse: Using the same address for multiple transactions make it easy to link payments to the same user [^17] This works because address are pseudonymous but not private.;
  • Temporal analysis: If transactions occur in rapid succession or at predictable intervals, they may be linked to the same entity;
@@
Existing countermeasures primarily aim to obscure input-output relations within transactions by using mixing protocols and preventing address reuse. Mixing protocols disrupt common-ownership heuristics and help mitigate change-address attacks. Address-changing mechanisms, such as generating a fresh sender or receiver address for each transaction, further reduce linkability. In addition, using network-level anonymity tools helps conceal user identities. Finally, the simplest defense against dust attacks is to ignore dust transactions, thereby preventing adversaries from linking addresses. A review of techniques for improving unlinkability follows.
Mixing Protocols
Different mixing protocols have been proposed (see Appendix A, Table LABEL:tab:sota-mixing). The naïve approach for mixing transactions consists in a centralized service that collects transactions, and publishes on the ledger a single large transaction with multiple inputs and outputs. This hides the information regarding which inputs flows towards which output, thus limiting (if not preventing) linkability between sender and receiver. The Mixcoin centralized protocol considers an accountable mixer, which offers a signed warranty that will enable his customers to unambiguously prove if the mix has misbehaved (Bonneau et al., 2014). Blindcoin (Valenta and Rowan, 2015) extends this protocol by considering blind signatures for the warrants; this hides the link between input and output addresses of a user even from the mixer, achieving full unlinkability. Tumblebit (Heilman et al., 2016) prevents coin thief using an on-chain escrow and ZKPs. In BlindHub (Qin et al., 2023), mixing is performed within payment channel hubs, where the hub acts as tumbler and uses blind signatures to hide the customers’ addresses from itself. Obscuro (Tran et al., 2018) is a mixing protocol that utilizes Trusted Execution Environments (TEEs) and enables participants to verify the mixer’s integrity via remote attestation. [^18] Obscuso is specifically designed to defend against two threats: the participation rejection attack, where malicious users try to manipulating mixing by withdrawing their deposits before the TEE reads them, and the blockchain forking attack, where adversaries present stale blocks to mislead the mixer. Numerous papers also investigate decentralized protocols. CoinJoin (Maxwell, 2013) is a technique that allows multiple users to combine their transactions into a single joint transaction; however, the author only describes the mixing idea. CoinShuffle (Ruffing et al., 2014) allow to combine multiple users’ inputs and outputs in a single transaction but may compromise anonymity if transaction values differ or the anonymity set is small. Later, Ruffing et al. (Ruffing et al., 2016) proposed CoinShuffle++, which uses a more efficient mechanism for anonymity, i.e., Dining Cryptographers Networks (DC-nets) instead of mix-nets: while a mix-net requires sequential processing, hence a number of communication rounds linear in the number of participants, DC-nets enable to process mixing in parallel, hence it requires only a constant number of communication rounds. Xim (Bissias et al., 2014) is a multi-round protocol for anonymously finding mix partners based on advertisements placed on the blockchain, with the property that no outside party can identify participants that pair up. SecureCoin (Ibrahim, 2017) creates an aggregated temporary deposit address and requires the collaboration of involved peers to generate novel destination addresses and collectively redistribute coins. Leveraging Confidential Transactions (which nonetheless is available only in a soft fork of Bitcoin), ValueShuffle (Ruffing and Moreno-Sanchez, 2017) provides full privacy besides parties anonymity. Confidentiality enables mixing funds of different values without compromising unlinkability. CoinParty (Ziegeldorf et al., 2018) is a decentralized mixing protocol that distributes the mixing process across a group of semi-trusted parties. They use an MPC protocol to jointly perform sensitive operations (i.e., encryption, decryption, and address shuffling), ensuring that no single party can link input and output addresses. Threshold signatures is used to enable claiming funds upon agreement of a majority of mixing nodes. It also includes mechanisms to detect and exclude misbehaving participants, increasing the overall protocol reliability.
Decentralized tumblers have also been designed using smart contracts on Ethereum. Möbius (Meiklejohn and Mercer, 2018) uses ring signatures and stealth addresses to obfuscate the senders and recipients associations. First, senders furnish both their funds and stealth keys to the smart contract; then, each recipient creates a ring signature to withdraw their funds from the contract and transfer them to an ephemeral address. The size of the anonymity set is limited to the size of the ring, and the gas cost of the withdrawing transaction increases linearly with the size of the ring. MixEth (Seres et al., 2019) aims to minimize gas costs by utilizing Neff’s verifiable shuffle (Neff, 2001) and off-chain messages. Once the sender deposits coins, the MixEth contract enters rounds of shuffling and challenging, during which anyone can challenge the correctness of the preceding shuffle. When recipients believe that sufficient shuffling has been performed, they can withdraw their funds from the MixEth contract.
Mixing protocols are still vulnerable to transaction graph analysis to some extent. For example, if the mixing process does not involve adequate randomization or if there are patterns in the amounts or timing of transactions, likely connections can be identified, even after several rounds of mixing. To maximize privacy, participants need to combine mixing protocols with other privacy-preserving techniques and follow best practices (e.g., distributing coins across multiple addresses) to reduce the risk of being traced through graph analysis. Further details on mixing solutions for Bitcoin and Ethereum, e.g., in (Arbabi et al., 2023), and on transaction graph analysis, e.g., in (Ron and Shamir, 2013; Fleder et al., 2015; Chan and Olmsted, 2017).
@@
Privacy-native solutions with public ledgers leverage a variety of strategies to improve unlinkability, focusing primarily on address-changing, transaction mixing, and use of ZKPs. Each of these techniques adds a layer of privacy, with trade-offs in terms of anonymity, computational complexity, and vulnerability to analysis. One-time destination addresses, stealth addresses (e.g., (Van Saberhagen, 2013; Noether et al., 2016)), or updatable public keys (e.g., (Papadoulis et al., 2024; Fauzi et al., 2019)) help obtain recipient unlinkability. While this ensures the recipient’s privacy, it does not hide transaction amounts. Protocols leveraging updatable public keys require the receiver to periodically scan the ledger, adding complexity. With ring signatures (e.g., (Noether et al., 2016; Van Saberhagen, 2013)), a sender creates a ring with multiple decoy inputs, making it impossible to determine which input belongs to the spender. However, the privacy protection is limited by the number of decoys in the ring, and sophisticated heuristics can sometimes reduce the anonymity set. For the most robust privacy, ZKPs offer a powerful solution (e.g., (Ben Sasson et al., 2014; Poelstra, 2016)). They allow transactions to prove their validity without revealing any details about the sender, receiver, or transaction amount. With this approach, transactions are recorded as cryptographic proofs (instead of sender-receiver addresses), ensuring that outputs cannot be linked to inputs. Among ZK-based approaches, we mention the commitments and nullifiers strategy (e.g., (Ben Sasson et al., 2014; Tomescu et al., 2022)): a coin contains a commitment to its serial number, and spending the coin involves publishing a nullifier, i.e., a pseudo-random value derived from the coin serial number. This ensures unlinkability to the original coin, while marking it as spent. However, designing a system with commitments and nullifiers is more complex than ring signatures or stealth addresses, and is also more computational demanding (see Section 4.5).
@@ 4.3.6. Plausible Deniability
With plausible deniability, we intend whether a user can deny intentional participation in a transaction, preventing others and adversaries from proving their involvement (e.g., (Almashaqbeh and Solomon, 2022; Ziegeldorf et al., 2018)). Existing works are roughly evenly divided between those supporting this property (e.g., (Noether et al., 2016; Poelstra, 2016; Narula et al., 2018; Kiayias et al., 2022; Tomescu et al., 2022; Wüst et al., 2022; Goodell et al., 2023; Papadoulis et al., 2024)) and those not (e.g., (Maxwell, 2013; Van Saberhagen, 2013; Garman et al., 2017; Bünz et al., 2020; Gross et al., 2021; Baudet et al., 2023; Guan et al., 2022; Yu et al., 2024)). Basically, protocols “on-demand”, such as mixing protocols, subsume an explicit user intention in using the additional service to obtain a privacy feature (such as unlinkability). Likewise, overlay systems implementing an anonymous crypto-asset systems, such as Zerocoin (Miers et al., 2013), Zerocash (Ben Sasson et al., 2014), EZC (Androulaki and Karame, 2014), Zether (Bünz et al., 2020), require explicit intention to convert a native asset to a privacy-preserving one. Conversely, systems that force anonymity or unlinkability as default, readily provide the property of plausible deniability. Among these systems, we mention Monero (Noether et al., 2016), MimbleWimble (Poelstra, 2016), UTT (Tomescu et al., 2022), PEReDi (Kiayias et al., 2022), or Platypus (Wüst et al., 2022), that do not provide the functionality of combining anonymous and public transactions.
@@ 4.4. Auditability
Auditability refers to the ability to examine, track, and verify transactions or balances while preserving a certain degree of privacy. It is a key feature of the third generation of privacy-preserving digital payment systems and, in particular, CBDCs. Two concepts closely related to auditability are: policy and accountability. A policy refers to the rules enforced by the protocol or by users to govern behaviors (e.g., to restrict who can transact and under what conditions). While policies involve ex-ante enforcement, auditability supports post-facto verification. Accountability is a more general concepts related to linking actions to individuals, e.g., through identity-binding mechanisms used by registration authorities to comply with AML/CFT regulations. Full privacy often conflicts with compliance: if transactions are completely opaque, regulators have no oversight (Ballaschk and Paulick, 2021). We identify the following approaches that enable or support auditability: anonymity budget, operating limits, revocable anonymity and user tracing, coin tracing, and verification without disclosure.
@@ 4.4.2. Operating Limits
Operating limits indicate constraints on transactions and accounts, such as maximum value, user balance, total sent/received funds, or transaction count within a time window. Compliance is generally proven in zero-knowledge. Garman et al. (Garman et al., 2017) extend Zerocash with auditability features, including regulatory closure (i.e., restricting transactions to assets of the same type), spending limits, selective user inquiries, and coin tracing. A central idea is the per-transaction spending limit: transactions above the limit are invalid unless signed by an authority. Another approach uses a per-user counter incremented with each outgoing transaction. The system can then reject transactions beyond a threshold. Gross et al. (Gross et al., 2021) build on this with a two-layer account-based system for implementing a CBDC, where each account tracks epoch turnover (total spent per epoch), proven with ZKPs. Similarly, AQQUA (Papadoulis et al., 2024) maintains multiple counters per account (balance, sent, received), publicly stored as three commitments. A registration authority links users identities to their (possibly, multiple) public keys, while an audit authority check compliance with five policy types: spending/receiving limits, transaction value limits, inactivity over a time period, selective transaction disclosure. During an audit, the authority selects a user by his public key and requests the opening of two historical account state snapshots (for every user’s accounts). After audits, the user randomizes his accounts (and provides a ZKP of valid update), thereby restoring his privacy. Other systems also rely on counters. Xue et al. (Xue et al., 2023) use per-user counters stored in a Merkle tree managed by regulators, which are also in charge of user on-boarding. ZKPs [^19] Specifically, zk-SNARKs are used for arithmetic statements and Sigma Protocols for algebraic statements. on counters enable policies on total amount and frequency of fund transfers within a time period. When suspicious transactions are detected, the participants’ identities can be recovered by regulators. Chen et al. (Chen et al., 2020) define policies on transaction limits, rates, and value, also verified with ZKPs.
@@ 4.4.3. Revocable Anonymity and User Tracing
Some schemes rely on special digital signatures that expose the sender on double spending (e.g., (Hayes, 1990; Zhang et al., 2019; Beer et al., 2024)). Others embed encrypt identifiers, keys, or metadata into tags, accessible only to tracing authorities (e.g., (Cecchetti et al., 2017; Garman et al., 2017; Zhang et al., 2019; Liu et al., 2022; Lin et al., 2020; Li et al., 2021; Kiayias et al., 2022)). This is often referred to as the trapdoor approach. Many systems also introduce a registration authority that links cryptographic keys to real-world identities (e.g., (Zhang et al., 2019; Liu et al., 2022)). For instance, Cecchetti et al. (Cecchetti et al., 2017) use oblivious RAM to manage transactions across banks, while Liu et al. (Liu et al., 2022) encrypt user identities under the central bank’s key. Zhang et al. (Zhang et al., 2019) rely on a supervision authority with decryption power. Garman et al. (Garman et al., 2017) propose a tracing mechanism where each user receives a unique encryption key: if the user is being traced, the key is a randomized version of the authority’s key; otherwise, it is derived from a null key. Androulaki et al. (Androulaki et al., 2020) encrypt transaction data under the auditors of both sender and receiver, while Xue et al. (Xue et al., 2023) scope tracing by encoding user identities with epoch identifiers. The epoch identifier enables temporal scoping, which limits the impact of key compromising and can enable granular tracing. Other systems derive one-time keys from long-term keys, enabling tracing by recovering the latter. Lin et al. (Lin et al., 2020) let users perform transactions using anonymous keys derived from their public addresses. A manager can use his key and the anonymous key to recover the user public address. Traceable Monero (Li et al., 2021) tags payee one-time keys: an authority can decrypt the tag and perform backward and forward tracing (i.e., identify the accounts that sent/receive funds to/from the current one).
A common concern is the trust concentration in a single tracing authority. To address this, Yu et al. (Yu et al., 2024) split tracing between two authorities via encrypted “marks” requiring joint computation, while PEReDi (Kiayias et al., 2022) (and Sarencheh et al. (Sarencheh et al., 2025)) distribute decryption among multiple parties, requiring quorum approval to identify users and access transaction metadata.

### v3

@@ 4.3.5. Unlinkability
  • Address reuse: Using the same address for multiple transactions makes it easy to link payments to the same user [^26] This works because address are pseudonymous but not private.;
  • Temporal analysis: If transactions occur in rapid succession or at predictable intervals, they may be linked to the same entity, e.g., due to automated systems that process transactions in batches or periodically;
@@
Existing countermeasures primarily aim to obscure input-output relations within transactions by using mixing protocols and preventing address reuse. Mixing protocols disrupt common-ownership heuristics and help mitigate change-address attacks (an analysis of mixing protocols in  Appendix A.3.2). Address-changing mechanisms, such as generating a fresh sender or receiver address for each transaction, further reduce linkability. In addition, using network-level anonymity tools helps conceal user identities. Finally, the simplest defense against dust attacks is to ignore dust transactions, thereby preventing adversaries from linking addresses.
@@
Privacy-native solutions with public ledgers leverage a variety of strategies to improve unlinkability, focusing primarily on address-changing, transaction mixing, and use of ZKPs. Each of these techniques adds a layer of privacy, with trade-offs in terms of anonymity, computational complexity, and vulnerability to analysis. One-time destination addresses, stealth addresses (e.g., (Van Saberhagen, 2013; Noether et al., 2016)), or updatable public keys (e.g., (Papadoulis et al., 2024; Fauzi et al., 2019)) help obtain recipient unlinkability. While this ensures the recipient’s privacy, it does not hide transaction amounts. With ring signatures (e.g., (Noether et al., 2016; Van Saberhagen, 2013)), the privacy protection is limited by the number of decoys in the ring, and sophisticated heuristics can sometimes reduce the anonymity set. For the most robust privacy, ZKPs offer a powerful solution (e.g., (Ben Sasson et al., 2014; Poelstra, 2016)). They allow transactions to prove their validity without revealing any details about the sender, receiver, or transaction amount. With this approach, transactions are recorded as cryptographic proofs (instead of sender-receiver addresses), ensuring that outputs cannot be linked to inputs. Among ZK-based approaches, we mention the commitments and nullifiers strategy (e.g., (Ben Sasson et al., 2014; Tomescu et al., 2022)): a coin contains a commitment to its serial number, and spending the coin involves publishing a nullifier, i.e., a pseudo-random value derived from the coin serial number. This ensures unlinkability to the original coin, while marking it as spent. However, designing a system with commitments and nullifiers is more complex than ring signatures or stealth addresses, and is also more computationally demanding (see Section 4.5).
@@ 4.3.6. Plausible Deniability
Plausible deniability indicates whether users can deny intentional participation in a transaction, preventing others from proving their involvement (e.g., (Almashaqbeh and Solomon, 2022; Ziegeldorf et al., 2018)). Existing works are roughly evenly divided between those supporting this property (e.g., (Noether et al., 2016; Poelstra, 2016; Narula et al., 2018; Kiayias et al., 2022; Tomescu et al., 2022; Wüst et al., 2022; Goodell et al., 2023; Papadoulis et al., 2024)) and those not (e.g., (Maxwell, 2013; Van Saberhagen, 2013; Garman et al., 2017; Bünz et al., 2020; Gross et al., 2021; Baudet et al., 2023; Guan et al., 2022; Yu et al., 2024)). Basically, protocols “on-demand”, such as mixing protocols, subsume an explicit user intention in using the additional service to obtain a privacy feature (such as unlinkability). Likewise, overlay systems implementing anonymous crypto-asset systems, such as Zerocoin (Miers et al., 2013), Zerocash (Ben Sasson et al., 2014), EZC (Androulaki and Karame, 2014), Zether (Bünz et al., 2020), require explicit intention to convert a native asset to a privacy-preserving one. Conversely, systems that force anonymity or unlinkability as default, readily provide the property of plausible deniability. Among these systems, we mention Monero (Noether et al., 2016), MimbleWimble (Poelstra, 2016), UTT (Tomescu et al., 2022), PEReDi (Kiayias et al., 2022), or Platypus (Wüst et al., 2022), that do not provide the functionality of combining anonymous and public transactions.
@@ 4.4. Auditability
While privacy is a core requirement of digital payment systems, complete opacity is often incompatible with regulatory and operational constraints. In regulated environments (e.g., CBDCs), some form of oversight is required to support AML/CFT controls, institutional accountability, and post-incident analysis (e.g., fraud, theft, protocol failures). Auditability should not be conflated with full transparency or continuous surveillance. Prior work has highlighted a broad design space between cash-like anonymity and full transaction observability, in which selective oversight can be enabled without pervasive disclosure (e.g., (Goodell and Aste, 2019; Pocher and Veneris, 2022; Ballaschk and Paulick, 2021; Michalopoulos et al., 2025)). In this sense, auditability enables authorized parties to verify compliance properties or selectively access transaction information, rather than making transactions universally visible. From a design perspective, auditability introduces a controlled relaxation of privacy guarantees. Increasing audit power typically strengthens trust assumptions or governance requirements and may reduce anonymity. Therefore, existing systems differ in what can be audited, when auditing may occur, and who is trusted to perform it. Following a regulation-by-design approach, auditability mechanisms can be embedded directly into the protocol to reconcile compliance with privacy. As discussed in Sections 4.1 and 4.2, the feasibility and scope of auditability mechanisms depend strongly on protocol architecture and on the underlying value representation. As such, auditability represents an explicit design concern of third generation systems, rather than an isolated feature.
Auditability, Policy, and Accountability
Auditability refers to the ability to examine, track, and verify transactions and balances while preserving a certain degree of privacy. It is a key feature of the third generation of privacy-preserving digital payment systems and, in particular, CBDCs. Auditability operates between these notions by supporting post-facto verification: it allows authorized parties to assess compliance or investigate suspicious behavior after transactions have occurred, and only within a prescribed scope. Two related concepts clarify its role: policy and accountability. A policy specifies admissible behavior and is typically enforced ex-ante by preventing invalid transactions. Accountability concerns the ability to associate actions with responsible entities, often via identity-binding mechanisms established at registration time.
What is audited, in practice?
Usually, auditability does not grant auditors unrestricted access to transaction data. Instead, it provides structured access to specific protocol state. This typically includes persistent state associated with users or tokens (e.g., counters or commitments), encrypted auxiliary metadata, and ZKPs attesting that policy constraints were respected. Importantly, auditability is rarely tied to a single account or address. To prevent trivial circumvention (e.g., by generating fresh accounts), audit-relevant state is anchored to identity credentials, token histories, or cryptographically bound counters that persist across transactions. As a result, address rotation does not reset spending history or policy enforcement. Auditability mechanisms represent constraints on system evolution rather than direct inspection tools: they regulate how hidden state may evolve over time and enable authorized parties to verify compliance without pervasive disclosure.
Mechanisms
We identify the following approaches that enable or support auditability: anonymity budget, operating limits, revocable anonymity and user tracing, coin tracing, and verification without disclosure.
@@ 4.4.2. Operating Limits
Operating limits indicate constraints on transactions and accounts, such as maximum value, user balance, total sent/received funds, or transaction count within a time window. Compliance is generally proven in zero-knowledge. Garman et al. (Garman et al., 2017) extend Zerocash with auditability features, including regulatory closure (i.e., restricting transactions to assets of the same type), spending limits, selective user inquiries, and coin tracing. A central idea is the per-transaction spending limit: transactions above the limit are invalid unless signed by an authority. Another approach uses a per-user counter incremented with each outgoing transaction. The system can then reject transactions beyond a threshold. Gross et al. (Gross et al., 2021) build on this with a two-layer account-based system for implementing a CBDC, where each account tracks epoch turnover (total spent per epoch), proven with ZKPs. Similarly, AQQUA (Papadoulis et al., 2024) maintains multiple counters per account (balance, sent, received), publicly stored as three commitments. A registration authority links users identities to their (possibly, multiple) public keys, while an audit authority check compliance with five policy types: spending/receiving limits, transaction value limits, inactivity over a time period, selective transaction disclosure. During an audit, the authority selects a user by his public key and requests the opening of two historical account state snapshots (for every user’s accounts). After audits, the user randomizes his accounts (and provides a ZKP of valid update), thereby restoring his privacy. Other systems also rely on counters. Xue et al. (Xue et al., 2023) use per-user counters stored in a Merkle tree managed by regulators, which are also in charge of user on-boarding. ZKPs [^27] Specifically, zk-SNARKs are used for arithmetic statements and Sigma Protocols for algebraic statements. on counters enable policies on total amount and frequency of fund transfers within a time period. When suspicious transactions are detected, the participants’ identities can be recovered by regulators. Chen et al. (Chen et al., 2020) define policies on transaction limits, rates, and value, also verified with ZKPs. Across these proposals, auditability is achieved by maintaining cumulative state (e.g., counters or commitments) whose evolution is constrained by policy.
@@ 4.4.3. Revocable Anonymity and User Tracing
Some schemes rely on special digital signatures that expose the sender on double spending (e.g., (Hayes, 1990; Zhang et al., 2019; Beer et al., 2024)). For example, Chaum-Fiat-Naor (Chaum et al., 1990) embed identifying information that is revealed upon double spending, while Schoenmakers (Schoenmakers, 1998) improves efficiency using “once-concealed, twice-revealed” proofs. Brands (Brands, 1994) replaces blind signatures with credential-based encodings to support selective disclosure, and Camenisch-Lysyanskaya (Camenisch and Lysyanskaya, 2001) signatures further generalize these ideas to multi-attribute tokens with flexible revocation policies.
Other approaches embed encrypted identifiers, keys, or metadata into tags, accessible only to tracing authorities (e.g., (Cecchetti et al., 2017; Garman et al., 2017; Zhang et al., 2019; Liu et al., 2022; Lin et al., 2020; Li et al., 2021; Kiayias et al., 2022)). This is often referred to as the trapdoor approach. Many systems also introduce a registration authority that links cryptographic keys to real-world identities (e.g., (Zhang et al., 2019; Liu et al., 2022)). For instance, Barki et al. (Barki and Gouget, 2020) adopt this model in a permissioned setting, embedding encrypted identity information into transactions and enabling a designated revocation authority to deanonymize users when required. Cecchetti et al. (Cecchetti et al., 2017) use oblivious RAM to manage transactions across banks, while Liu et al. (Liu et al., 2022) encrypt user identities under the central bank’s key. Zhang et al. (Zhang et al., 2019) rely on a supervision authority with decryption power. Li et al. (Li et al., 2019) propose an identity tag encrypted under a tracing authority’s key. Garman et al. (Garman et al., 2017) propose a tracing mechanism where each user receives a unique encryption key: if the user is being traced, the key is a randomized version of the authority’s key; otherwise, it is derived from a null key. Androulaki et al. (Androulaki et al., 2020) encrypt transaction data under the auditors of both sender and receiver, while Xue et al. (Xue et al., 2023) scope tracing by encoding user identities with epoch identifiers. The epoch identifier enables temporal scoping, which limits the impact of key compromising and can enable granular tracing (e.g., to investigate suspicious behavior in a particular time window). Other systems derive one-time keys from long-term keys, enabling tracing by recovering the latter. Lin et al. (Lin et al., 2020) let users perform transactions using anonymous keys derived from their public addresses. A manager can use his key and the anonymous key to recover the user’s public address. Traceable Monero (Li et al., 2021) tags payee one-time keys: an authority can decrypt the tag and perform backward and forward tracing (i.e., identify the accounts that sent/receive funds to/from the current one).
A common concern is the trust concentration in a single tracing authority. To address this, Yu et al. (Yu et al., 2024) split tracing between two authorities via encrypted “marks” requiring joint computation, while PEReDi (Kiayias et al., 2022) (and Sarencheh et al. (Sarencheh et al., 2025)) distribute decryption among multiple parties, requiring quorum approval to identify users and access transaction metadata. Similarly, PICTURE (Zhang et al., 2024) adopts a threshold-based regulation model, where unlinkable transactions embed encrypted trace records that can be opened only by a quorum of authorities to recover sender and receiver identities.